# On TinyML and Cybersecurity: Electric Vehicle Charging Infrastructure Use Case

### v2

@@ I Introduction
AIDS Anomaly-based Intrusion Detection Systems
@@
IDS Intrusion Detection Systems
@@
SIDS Signature-based Intrusion Detection Systems
@@
The implementation of TinyML plays a pivotal role for improving cybersecurity in the domain of EVCI [25]. EVCIs are susceptible to various cybersecurity vulnerabilities due to their interconnected nature and the sensitive data they handle, such as payment information and real-time energy usage statistics. Integrating the deployment of TinyML algorithms in EV aims to establish lightweight yet robust security measures at the edge. This approach focuses on enhancing data privacy, reducing latency, and improving the resilience of critical charging infrastructures against potential cyber threats. As protecting EVs from cyber threats is essential, this paper includes a case study on using TinyML in EVCI cybersecurity.
@@
  • An examination of TinyML implementation on the ESP32 microcontroller, including detailed programming practices, communication protocols, and performance insights.
The paper is structured as follows: Section II serves as a background, introducing TinyML, discussing various cybersecurity threats, and highlighting TinyML’s benefits. Section III delves into challenges and constraints of using TinyML on devices with limited resources. Section IV elaborates on potential solutions to these challenges. Section V provides an overview of key tools and libraries that facilitate TinyML implementation. Section VI presents a comprehensive case study on the application of TinyML in EVCIs, focusing on the framework, communication protocols, and attack scenarios. Section VII presents the experimental setup and a detailed analysis of the results. Section VIII discusses the practical implementation on an ESP32 microcontroller. Section IX outlines future research directions. Finally, Section X concludes the paper by summarizing the key findings and contributions and highlighting the impact of TinyML on IoT and cybersecurity. Additionally, for a comprehensive list of acronyms used in this paper, please refer to Table I.
@@ II-A3 TinyML Applications
Wearable technology, for instance, highlights TinyML’s role in enhancing health, safety, and communication. Applications extend to condition monitoring in mobile machinery, where low-cost TinyML sensors offer solutions beyond expensive equipment [22]. In the field of healthcare, TinyML aids in improvements in hearing aid technology by distinguishing desired sounds in noisy environments. It also assists in health diagnosis, such as identifying COVID-19 symptoms [13].
@@
TinyML also plays a critical role in urban air quality monitoring by integrating into embedded systems to estimate CO2 emissions from vehicles using soft-sensor techniques. This approach enhances accuracy without relying on cloud-based servers [36].
@@
Furthermore, TinyML’s applications in ecological conservation, autonomous vehicles, and Brain-Computer Interfaces highlight its versatility. It aids in tasks such as tumor detection, emotional intelligence, smart navigation, and monitoring wildlife, such as sea turtles [13].
From recognizing sign language and handwriting to identifying medical face masks, interpreting gestures, and transcribing speech, the versatility of TinyML is evident [26]. These applications, along with many others, underline the broad potential of this technology across various industries and domains. The summary of TinyML applications is presented in Table III.
II-B Cybersecurity Threats
  • DoS and DDoS (Distributed Denial-of-Service) Attacks: These attacks aim to disable or restrict network services by overwhelming the system with excessive traffic or exploiting system weaknesses, which makes the service inaccessible to legitimate users [39].
  • Port Scan: Attackers probe network hosts to gather details about port numbers, operating systems, and applications by sending specific data and then analyzing the host’s response [40].
  • Botnet: A botnet is a network of malware-infected hosts controlled by malicious actors. These hosts, known as bots, are typically managed by a Command and Control (C&C) server, which can organize distributed attacks on other systems [41].
  • Infiltration: Infiltration refers to unauthorized actions on network hardware or software systems. The primary objective of Intrusion Detection Systems (IDS) is to detect and prevent these unauthorized actions [42].
  • Web Attacks: Including Brute Force, Cross-Site Scripting (XSS), and SQL Injection, these attacks target vulnerabilities in web services to gain unauthorized access or compromise data integrity [43, 44, 45].
  • Heartbleed: This attack undermines the encryption mechanism intended to secure online communications. Servers using a buggy version of Open Secure Socket Layer (OpenSSL) allow attackers to read the memory of systems, leading to the exposure of sensitive data such as private keys, user passwords, and confidential information [46].
  • Phishing and Spear Phishing: Phishing is a cyberattack strategy that deceives people into sharing sensitive data by using fraudulent websites, emails, and social engineering tricks. This leads victims to reveal personal, financial, or confidential information, often resulting in harmful consequences such as financial losses or data breaches [47, 48].
  • Ransomware: Malware that carries out malicious actions such as blocking access to personal data until a ransom is paid. This ransom is often demanded in cryptocurrency such as Bitcoin, making it challenging to trace transactions and evading law enforcement [49].
  • Man-In-The-Middle (MITM): Attacks intercepting communications between two parties to steal or manipulate data [50].
  • Zero-Day Attacks: A novel cyber threat targeting vulnerabilities unknown to the public and the cybersecurity community. Traditional signature-based detection methods often fail against these attacks, as they depend on known attack signatures, which zero-day threats lack. However, ML can effectively detect these attacks by analyzing statistical patterns in data. This approach, not dependent on known attack signatures, is better suited for combating novel threats [51].
@@ II-B Cybersecurity Threats
Infiltration Unauthorized actions on network hardware or software systems, detected by IDS [42].
@@ II-C The Role of TinyML in Enhancing Cybersecurity
As the IoT systems grow, the complexity of cyber threats increases, thus amplifying the importance of mitigating these threats. Unauthorized activities that threaten the confidentiality, integrity, and availability of these systems, known as intrusions, must be counteracted [52]. For this purpose, IDS are crucial for detecting unauthorized access, misuse, and breaches in computer networks and systems. They operate by monitoring and analyzing network traffic or system activities for malicious patterns or anomalies that indicate potential security threats [53]. There are two different types of IDSs: Signature-based Intrusion Detection Systems (SIDS) and Anomaly-based Intrusion Detection Systems (AIDS) [52].
SIDS operate by matching patterns to identify known attacks, comparing intrusion signatures against a pre-existing database [52]. Incorporating supervised ML algorithms in SIDS enhances their capability to recognize known threats. The key limitation here is that supervised learning models are trained on labeled datasets and face challenges with new, unfamiliar threats. Maintaining and processing a large database for SIDS may require considerable computational resources [54]. Additionally, the centralized structure, common in traditional cybersecurity systems such as SIDS, may introduce certain vulnerabilities, including single points of failure and privacy issues [55].
In contrast, AIDS do not rely solely on known patterns. Instead, this approach uses ML and statistical methods to establish a baseline of normal behavior for a system. Deviations from this baseline are flagged as potential intrusions [52]. ML models, especially unsupervised learning models, are capable of autonomously identifying patterns and anomalies, independent of predefined threat criteria. These models contribute to the creation of sophisticated frameworks for detecting malware, classifying spam, and identifying intrusions and anomalies. To facilitate this process, powerful servers with extensive computational resources are used to analyze large datasets and identify vulnerabilities [9].
Furthermore, AIDS, often referred to as behavior-based IDS, exhibit a high False Alarm Rate (FAR). This problem can lead to unnecessary panic, wasted resources, and potential system downtime while verifying the authenticity of the alert. Additionally, the increasing prevalence of zero-day attacks, which exploit unknown vulnerabilities, poses a significant challenge to the effectiveness of AIDS. These attacks require the IDS to swiftly adapt, a task that is particularly challenging due to the inherent nature of these unforeseen threats [54].
@@ III-A3 Computational Resources
TABLE VI: Summary of TinyML Challenges
Category Challenges
@@ III-C4 Real-time Inflexibility
The summary of TinyML challenges can be found in Table VI.
IV Potential Solutions to TinyML Challenges
@@ V TinyML Tools and Libraries
  • Model Optimization and Conversion Tools:
@@
  • Development Environments and Frameworks:
@@
  • Specialized Frameworks and Libraries:
@@
  • AI Integration and Industry-Specific Tools:
@@
VI Case Study: Electric Vehicle Charging Infrastructure
@@
It is important to note that EVCSs are not without their challenges, as they are exposed to various threats and potential cyber attacks [25]. These attacks can manifest in several ways, including tampering with charging processes through WiFi connections, bypassing authentication mechanisms, injecting malicious software, and disabling chargers [99].
Given potential cyber threats to EVCSs, traditional ML models offer a viable defense [58]. However, considering the limitations of these traditional models, such as their resource constraints on IoT devices, TinyML emerges as a more suitable alternative. TinyML stands out due to its inherent advantages in security and privacy, real-time processing capabilities, and cost-effectiveness which makes it an optimal choice for enhancing the safety and efficiency of EVCIs.
VI-A Framework and Standards
@@ VI-A3 SAE
VI-A4 CHAdeMO
CHAdeMO standard encompasses both physical and communication aspects. For the physical aspect, CHAdeMO defines the physical connector and charging specifications for high-voltage DC fast charging. The CHAdeMO protocol is used for communication between the EVCS and the EV, ensuring optimal and fast charging based on the EV’s commands. CHAdeMO chargers are popular in Japan [101] and Europe [103]. CHAdeMO, while enabling Vehicle-to-Grid (V2G) operations, has certain limitations that raise concerns. From the communication side, one of the main drawbacks is the lack of secure communication features, relying instead on Controller Area Network (CAN) communication. CAN enables electronic components within a vehicle to communicate with each other without relying on a central computer. This exposes the vehicle’s CAN bus to potential security risks, as unencrypted communication leaves it vulnerable to control or programming by malicious entities. Although CHAdeMO follows IEC standards for charging and digital communication, it currently lacks robust security measures. There is recognition of the importance of security, and efforts are being made to develop a unified communication protocol that would utilize Ethernet, TCP/IP, PKI infrastructure, and TLS encryption for enhanced security, similar to how TLS is employed on the Internet  [107].
VI-A5 IEC
@@
VI-B Communication Frameworks for EV Charging
@@
VI-B1 IoT-Enabled Smart Charging Network
@@
VI-B2 Mobile Applications within the Charging Ecosystem
@@
VI-B3 Power and Communication Flows in Electric Vehicle Charging Infrastructure
@@
VI-C Attack Scenario
The EVCI is increasingly vulnerable to a wide range of cyber-attacks, highlighting the critical need for robust security measures. Cyber-attacks on EVCI can be classified into local and remote attacks, taking advantage of the attacker’s physical proximity to the infrastructure or targeting the system via digital communication channels, respectively. Local attacks often involve the manipulation of electromagnetic fields to inject or alter data in sensors or relays, compromising the integrity of the charging process. Conversely, remote attacks typically aim at the cloud-based components of the infrastructure, such as user access controls and payment gateways, seeking to steal sensitive data or disrupt operations by altering control signals. These vulnerabilities are not only confined to the physical hardware but also extend to the communication protocols and cloud services integral to EVCI operations [109].

### v3

@@ I Introduction
AIDS Anomaly-based Intrusion Detection System
@@
CCLCS Coordinated Charging Load Control System
@@
IDS Intrusion Detection System
@@
PTQ post-training quantization
QAT Quantization-Aware Training
@@
SIDS Signature-based Intrusion Detection System
@@
This paper aims to provide a comprehensive review of TinyML challenges and potential solutions in general cybersecurity contexts. Given that EVCIs are among the most representative and impactful use cases for TinyML [25, 26, 27, 28], we include a detailed case study on EVCI. This case study serves to illustrate how TinyML can be effectively applied to real-world IoT systems, offering a clear understanding of the practical applications and benefits of TinyML in enhancing cybersecurity.
The implementation of TinyML plays a pivotal role in improving cybersecurity in the domain of EVCI [27]. EVCIs are susceptible to various cybersecurity vulnerabilities due to their interconnected nature and the sensitive data they handle, such as payment information and real-time energy usage statistics. Integrating the deployment of TinyML algorithms in Electric Vehicles (EVs) aims to establish lightweight yet robust security measures at the edge. This approach focuses on enhancing data privacy, reducing latency, and improving the resilience of critical charging infrastructures against potential cyber threats. To illustrate these benefits, this paper includes a case study on using TinyML in EVCI cybersecurity.
@@
  • An experimental setup and practical implementation of TinyML on the ESP32 microcontroller, including detailed programming practices, communication protocols, and performance insights.
The paper is structured as follows: Section II serves as a background, introducing TinyML, discussing various cybersecurity threats, and highlighting TinyML’s benefits. Section III delves into the challenges and constraints of using TinyML on devices with limited resources. Section IV elaborates on potential solutions to these challenges, along with key tools and libraries that facilitate TinyML implementation. Section V presents a comprehensive case study on the application of TinyML in EVCIs, focusing on the framework, communication protocols, attack scenarios, and countermeasures. Section VI presents the experimental setup and a detailed analysis of the results. Section VII discusses the practical implementation of TinyML on an ESP32 microcontroller. Section VIII details future research directions. Finally, Section IX concludes the paper by summarizing the key findings and contributions and highlighting the impact of TinyML on IoT and cybersecurity. Additionally, for a comprehensive list of acronyms used in this paper, please refer to Table I.
@@ II-A3 TinyML Applications
  • Wearable technology highlights TinyML’s role in enhancing health, safety, and communication. Applications extend to condition monitoring in mobile machinery, where low-cost TinyML sensors offer solutions beyond expensive equipment [22].
  • In healthcare, TinyML aids in improvements in hearing aid technology by distinguishing desired sounds in noisy environments. It also assists in health diagnosis, such as identifying COVID-19 symptoms [13].
@@
  • TinyML plays a critical role in urban air quality monitoring by integrating into embedded systems to estimate CO2 emissions from vehicles using soft-sensor techniques. This approach enhances accuracy without relying on cloud-based servers [39].
@@
  • TinyML’s applications in ecological conservation, autonomous vehicles, and Brain-Computer Interfaces highlight its versatility. It aids in tasks such as tumor detection, emotional intelligence, smart navigation, and monitoring wildlife, such as sea turtles [13].
  • TinyML is also used for recognizing sign language and handwriting, identifying medical face masks, interpreting gestures, and transcribing speech [29].
The summary of TinyML applications is presented in Table III.
@@
Infiltration Unauthorized actions on network hardware or software systems, detected by IDSs [45].
@@
II-B Cybersecurity Threats
II-B1 DoS and DDoS (Distributed Denial-of-Service) Attacks
These attacks aim to disable or restrict network services by overwhelming the system with excessive traffic or exploiting system weaknesses, which makes the service inaccessible to legitimate users [42].
II-B2 Port Scan
Attackers probe network hosts to gather details about port numbers, operating systems, and applications by sending specific data and then analyzing the host’s response [43].
II-B3 Botnet
A botnet is a network of malware-infected hosts controlled by malicious actors. These hosts, known as bots, are typically managed by a Command and Control (C&C) server, which can organize distributed attacks on other systems [44].
II-B4 Infiltration
Infiltration refers to unauthorized actions on network hardware or software systems. The primary objective of Intrusion Detection System (IDS) is to detect and prevent these unauthorized actions [45].
II-B5 Web Attacks
Including Brute Force, Cross-Site Scripting (XSS), and SQL Injection, these attacks target vulnerabilities in web services to gain unauthorized access or compromise data integrity [46, 47, 48].
II-B6 Heartbleed
This attack undermines the encryption mechanism intended to secure online communications. Servers using a buggy version of Open Secure Socket Layer (OpenSSL) allow attackers to read the memory of systems, leading to the exposure of sensitive data such as private keys, user passwords, and confidential information [49].
II-B7 Phishing and Spear Phishing
Phishing is a cyberattack strategy that deceives people into sharing sensitive data by using fraudulent websites, emails, and social engineering tricks. This leads victims to reveal personal, financial, or confidential information, often resulting in harmful consequences such as financial losses or data breaches [54, 50].
II-B8 Ransomware
Malware that carries out malicious actions such as blocking access to personal data until a ransom is paid. This ransom is often demanded in cryptocurrency such as Bitcoin, making it challenging to trace transactions and evading law enforcement [51].
II-B9 Man-in-the-Middle (MitM)
Attacks intercepting communications between two parties to steal or manipulate data [52].
II-B10 Zero-Day Attacks
A novel cyber threat targeting vulnerabilities unknown to the public and the cybersecurity community. Traditional signature-based detection methods often fail against these attacks, as they depend on known attack signatures, which zero-day threats lack. However, ML can effectively detect these attacks by analyzing statistical patterns in data. This approach, not dependent on known attack signatures, is better suited for combating novel threats [53].
@@ II-C The Role of TinyML in Enhancing Cybersecurity
As the IoT systems grow, the complexity of cyber threats increases, thus amplifying the importance of mitigating these threats. Unauthorized activities that threaten the confidentiality, integrity, and availability of these systems, known as intrusions, must be counteracted [55]. For this purpose, IDSs are crucial for detecting unauthorized access, misuse, and breaches in computer networks and systems. They operate by monitoring and analyzing network traffic or system activities for malicious patterns or anomalies that indicate potential security threats [56]. There are two different types of IDSs: Signature-based Intrusion Detection System (SIDS) and Anomaly-based Intrusion Detection System (AIDS) [55].
SIDSs operate by matching patterns to identify known attacks, comparing intrusion signatures against a pre-existing database [55]. Incorporating supervised ML algorithms in SIDS enhances their capability to recognize known threats. The key limitation here is that supervised learning models are trained on labeled datasets and face challenges with new, unfamiliar threats. Maintaining and processing a large database for SIDS may require considerable computational resources [57]. Additionally, the centralized structure, common in traditional cybersecurity systems such as SIDS, may introduce certain vulnerabilities, including single points of failure and privacy issues [58].
In contrast, AIDSs do not rely solely on known patterns. Instead, this approach uses ML and statistical methods to establish a baseline of normal behavior for a system. Deviations from this baseline are flagged as potential intrusions [55]. ML models, especially unsupervised learning models, are capable of autonomously identifying patterns and anomalies, independent of predefined threat criteria. These models contribute to the creation of sophisticated frameworks for detecting malware, classifying spam, and identifying intrusions and anomalies. To facilitate this process, powerful servers with extensive computational resources are used to analyze large datasets and identify vulnerabilities [9].
Furthermore, AIDSs, often referred to as behavior-based IDSs, exhibit a high False Alarm Rate (FAR). This problem can lead to unnecessary panic, wasted resources, and potential system downtime while verifying the authenticity of the alert. Additionally, the increasing prevalence of zero-day attacks, which exploit unknown vulnerabilities, poses a significant challenge to the effectiveness of AIDS. These attacks require the IDS to swiftly adapt, a task that is particularly challenging due to the inherent nature of these unforeseen threats [57].
@@ III-B2 Limited Datasets
TABLE VI: Summary of TinyML Challenges
Category Challenges
@@ III-C4 Real-time Inflexibility
The summary of TinyML challenges can be found in Table VI. Addressing these challenges is crucial for the effective deployment of TinyML in EVCI. For example, EVs aim to minimize power consumption to extend battery life and enhance sustainable development. By optimizing TinyML solutions to operate within the stringent resource constraints of EVCI, we can enhance the security, efficiency, and reliability of these critical systems, ensuring they can meet the demands of future smart cities and the increasing adoption of EVs.
IV Potential Solutions, Tools, and Libraries for TinyML
@@ IV-G Tools and Libraries
IV-G1 Model Optimization and Conversion Tools
@@
IV-G2 Development Environments and Frameworks
@@
IV-G3 Specialized Frameworks and Libraries
@@
IV-G4 AI Integration and Industry-Specific Tools
@@
V Case Study: Electric Vehicle Charging Infrastructure
Among the numerous applications of TinyML across various sectors, EVCIs stand out as a particularly crucial and challenging domain. As the world moves towards smarter and more connected urban environments, the role of EVCI becomes increasingly significant in ensuring the seamless integration of electric vehicles into the urban landscape.
@@ V Case Study: Electric Vehicle Charging Infrastructure
However, the expansion of EVCSs introduces significant cybersecurity challenges. These systems are vulnerable to various threats, including tampering with charging processes through WiFi connections, bypassing authentication mechanisms, injecting malicious software, and disabling chargers [102]. Given these potential cyber threats, there is an urgent need for robust security solutions.
While traditional ML models provide a baseline defense against cyber threats in EVCSs, they often face resource constraints when deployed on IoT devices, which is a critical consideration in the expansion of EVCI [61]. In this context, TinyML emerges as a more suitable alternative. TinyML offers substantial advantages in terms of cost and energy effectiveness, while still providing enhanced security and privacy, as well as real-time processing capabilities, making it an optimal choice for improving the safety and efficiency of EVCIs.
V-A Framework and Standards
@@ V-A3 SAE
V-A4 CHAdeMO
CHAdeMO standard encompasses both physical and communication aspects. For the physical aspect, CHAdeMO defines the physical connector and charging specifications for high-voltage DC fast charging. The CHAdeMO protocol is used for communication between the EVCS and the EV, ensuring optimal and fast charging based on the EV’s commands. CHAdeMO chargers are popular in Japan [103] and Europe [105]. CHAdeMO, while enabling Vehicle-to-Grid (V2G) operations, has certain limitations that raise concerns. From the communication side, one of the main drawbacks is the lack of secure communication features, relying instead on Controller Area Network (CAN) communication. CAN enables electronic components within a vehicle to communicate with each other without relying on a central computer. This exposes the vehicle’s CAN bus to potential security risks, as unencrypted communication leaves it vulnerable to control or programming by malicious entities. Although CHAdeMO follows IEC standards for charging and digital communication, it currently lacks robust security measures. There is recognition of the importance of security, and efforts are being made to develop a unified communication protocol that would utilize Ethernet, Transmission Control Protocol/Internet Protocol (TCP/IP), PKI infrastructure, and TLS encryption for enhanced security, similar to how TLS is employed on the Internet  [26].
V-A5 IEC
@@
V-B Communication Frameworks for EV Charging
@@
V-B1 IoT-Enabled Smart Charging Network
@@
V-B2 Mobile Applications within the Charging Ecosystem
@@
V-B3 Power and Communication Flows in Electric Vehicle Charging Infrastructure
@@
V-C Cyber Attack Scenarios in EVCI
EVCSs provide power to EVs through an IoT-enabled infrastructure operating on specialized firmware. However, an attacker can disrupt the charging process if they gain access to the charging equipment via a Wi-Fi connection [110]. Protecting EVCI from cyber threats is critical, as a wide range of potential attacks—including DoS, firmware manipulation, and malicious code injection—could disrupt charging operations, compromise user privacy, or destabilize the electric grid. Addressing these risks requires a comprehensive cybersecurity strategy that includes intrusion detection, standardized security practices, and robust authentication mechanisms [111].
Cybersecurity attacks on EVCI can manifest in various forms. Since the communication system serves as the backbone of EVCI, handling crucial operations such as EV scheduling, authentication, and grid integration, any compromise in this communication can leave the entire EVCI vulnerable [28]. Different motivations exist for attackers, ranging from electricity theft to disrupting a network of charging stations [112].
The interconnected nature of EVs and charging stations presents significant challenges in terms of security and reliability. EVCIs are exposed to a wide range of cybersecurity threats, such as DoS attacks, MitM attacks, eavesdropping, spoofing, False Data Injection (FDI), hacking, ransomware, Trojan viruses, malware injection, and data breaches. These threats can lead to personal or financial data breaches and hijacking of power or charging infrastructure [110, 113, 114].
Constant communication between EVSE and smart grids means that weaknesses in charging stations can have far-reaching consequences for the entire grid. The critical exchange of information occurs during vehicle charging or when users interact with their vehicles via software, creating a potential bottleneck where sensitive data and user privacy could be compromised. Securing this interconnected infrastructure is crucial for maintaining user privacy and ensuring reliable grid operations [113].
At the high-power, extreme-fast charging level, vulnerabilities are even more severe due to significant power flow. Attacks at this level can cause a loss of EV charging availability or safety concerns due to battery overcharging. Additionally, attackers could manipulate charging rates, block charging sessions, or even drain batteries, posing significant risks to both EV owners and charging station operators [114].
One prominent attack is the DoS, where an attacker overwhelms the CMS or EVSE with excessive traffic or high charging requests, causing system resources to crash or become unresponsive [112, 28, 110]. DoS attacks can be initiated in various ways, such as through botnets or flooding. In research by Manoj and Mohd, hackers used spoofed IPs of legitimate EVSEs to launch DoS attacks [28]. Another potential form of a DoS attack involves flooding the CAN bus of an EV with high volumes of traffic or spoofed messages, which could saturate the bus and disrupt the vehicle’s ability to perform normal operations, thereby preventing the vehicle from processing legitimate control messages. Additionally, if the communication link between the EV and the charging stations is flooded with excessive connection requests or malformed packets, the vehicle’s communication capabilities can be severely disrupted.
Gottumukkala et al.[112] mention that a DoS attack on an EVCS often begins with a spoofing attack, where an attacker compromises a device’s unique identifier, such as a MAC address, allowing them to masquerade as a legitimate user. Once the attacker gains unauthorized access to the system, they target the CMS or EVSE by flooding the network with excessive traffic or high charging requests. Attack variants like UDP or TCP/IP floods, low-rate DoS, ping floods, or Internet Control Message Protocol (ICMP) floods can take down a charging station or other nodes within the charging station network. This attack overwhelms system resources, causing the CMS to crash or become unresponsive, effectively preventing legitimate users from accessing charging services. To mitigate such attacks, EVCS operators should implement robust authentication mechanisms, traffic filtering, rate limiting, and anomaly detection techniques to identify and prevent abnormal traffic patterns[112].
Firmware manipulation is another significant threat, where attackers modify the charging station’s firmware to inject malicious code, potentially affecting the processes, control actions, and operating conditions of the system. This malicious code can take the form of viruses, ransomware, or SQL and XSS injection attacks [28]. Additionally, attackers can exploit EV connectors to install malware or manipulate charging settings, providing unauthorized access to the EVCS [110].
A lack of robust security in software modules makes charging stations and the cloud vulnerable to malware, enabling coordinated attacks that could shut down an entire network or impact the power grid by simultaneously activating multiple charging stations [112]. A more sophisticated attack, known as the Malicious Mode Attack (MMA), proposed by Weidong et al., targets the Coordinated Charging Load Control System (CCLCS) by sending forced oscillation control commands to multiple charging stations. This attack leads to high-amplitude forced oscillations in the power system by exploiting the weakly damped wide-area electromechanical mode, resulting in system instability and disrupting the power supply capacity of EVCSs [115].
The OCPP, one of the most widely used protocols in charging stations, lacks built-in security measures against cyberattacks, potentially providing unauthorized users with an access point and leaving vulnerabilities for MitM attacks. An intruder can exploit the minimum status duration function in OCPP to access the communication interface between the EVCS and the CMS. Once inside, the attacker manipulates EVs and charge points, altering basic start or stop transaction functions to overcharge batteries or prevent them from charging altogether, potentially creating hazardous situations or inconveniencing consumers who need to charge later [116].
MitM attacks are common in radio-based communications between EVSEs, Plug-In Electrical Vehicles (PEVs), and Building Energy Management Systems (BEMS). An attacker can corrupt data or take full control of a node, relaying incorrect information such as false charging station status [112]. Moreover, MitM attacks can occur during the authentication and authorization process between PEVs and charging stations or throughout the charging session. Despite following the ISO 15118 security standard, the communication channel is still susceptible to MitM attacks. In these attacks, an intruder intercepts and manipulates the data exchanged between the charging station and the vehicle. By masquerading as a legitimate participant, the attacker can gain unauthorized access, potentially resulting in vehicle, information, or energy theft [113].